PHOTOION-PAIR PRODUCTION IN HF, DF AND (HF)_{2}

\underline{QICHI HU}, JOHN W. HEPBURN, Department of Chemistry, University of British Columbia, Vancouver, BC V6T 1Z1, Canada.

The dissociation processes of HF/DF + hν → H^{+}/D^{+} + F^{−} and (HF)_{2} + hν → HF H^{+} + F^{−} have been studied through single photon excitation with coherent vacuum ultraviolet light. For HF and DF, the laser was scanned around the ion-pair dissociation energies. The spectra demonstrate that the ion-pair production could be greatly enhanced through the coupling between the ion-pair states and the Rydberg states converging to certain levels of HF^{+}/DF^{+}. The results from our current high resolution study lead to different conclusions than those resulting from lower resolution synchrotron experiments^{a}. For the ion-pair production of (HF)_{2}, some preliminary results and possible explanations will be presented.

aA. J. Yencha, A. Hopkirk, J. R. Grover, B-M. Cheng, H. Lefebvre-Brion, and F. Keller, J. Chem. Phys. 103, 2882 (1995)